A sporting car starting from rest accelerates 40 km/hr^2 for 30 min after which it traveled with a constant velocity of 1 hr. When the brakes where applied, it slow down at 2km/hr^2 until it stops. Find the total distance covered.
Answer by ikleyn(52781)

.
A sporting car starting from rest accelerates 40 km/hr^2 for 30 min after which it traveled with a constant velocity of 1 hr.
When the brakes where applied, it slow down at 2km/hr^2 until it stops. Find the total distance covered.
~~~~~~~~~~~~~~~~~~~~


By accelerating at 40 km/hr^2 for 30 minutes, the car will reach a velocity of


    v = a*t = 40 * 0.5 = 20 kilometers per hour.


It is not realistic for the sport car.


Therefore, I will stop my solution at this point, by saving my time for more reasonable purposes.